OpenStudy (anonymous):

Jimmy is making a rectangular garden. He wants it to be( x - 11 ) long and (x-1) yards wide. He needs a walkway 3 yards wide around the garden. If the total area of both garden and the walkway is 119 sq yards. solve for x. please explain.

OpenStudy (anonymous):

the given length is (x-11) and the width is (x-1), since there is a 3 yards extension for the pathway therefor both sides will be added by 3 given the area of both garden and pathway is 119. therefor the sides will be now length= (x-8) and width=(x+2)

OpenStudy (anonymous):

substitute to the formula  given the area find x. .

OpenStudy (anonymous):

can you do it now?

OpenStudy (anonymous):

(x-8)(x+2)=119 Is it 15 yds

OpenStudy (anonymous):

if that's the result of your computation using the formula and the given, then your answer is correct. . .

OpenStudy (anonymous):

Is your answer also 15 yds

OpenStudy (anonymous):

Thanks a lot for helping me to solve the problem !